### v1

@@ Abstract
We consider moving boundary problems for biophysics and introduce a new computational framework to handle the complexity of the bulk-surface PDEs. In our framework, interpretability is maintained by adapting the fast, generalizable and accurate structure preservation scheme in [29]. We show that mesh distortion is mitigated by adopting the pioneering work of [36], which is tied to an Arbitrary Lagrangian Eulerian (ALE) framework. We test our algorithms accuracy on moving surfaces with boundary for the following PDEs: advection-diffusion-reaction equations, phase-field models of Cahn-Hilliard type, and Helfrich energy gradient flows. We performed convergence studies for all the schemes introduced to demonstrate accuracy. We use a staggered approach to achieve coupling and further verify the convergence of this coupling using numerical experiments. Finally, we demonstrate broad applicability of our work by simulating state-of-the-art tests of biophysical models that involve membrane deformation.
@@
Biophysics, moving boundary problems, bulk-surface coupling, structure preservation, mesh redistribution, multiphysics coupling
{MSCcodes}
35R37, 74K15, 65N30, 76T99.
@@ 1 Introduction
Many biophysical phenomena in cell biology can be characterized by the feedback between form and function. Cells routinely change their shape during cell motility, wound healing, and development. Such shape changes are a result of coordination of multiple biochemical and mechanical processes within the cell. Not surprisingly, these phenomena have been studied extensively using models based on partial differential equations (PDEs) with different levels of biophysical complexity [78]. In concert with experiments, such models have provided insight into the mechanisms underlying the change in cell shape [31, 95, 63]. However, an open challenge in the field of computational biophysics is the development of robust computational tools for numerical solutions of the governing PDEs. For example, cellular membranes are mechanical boundaries that are characterized by lateral dimensions that are large compared to their thickness, appearing in a variety of structures in cells [99]. It has been proven that the properties of these surfaces are crucial for the genesis and maintenance of cellular membrane systems [16, 86]. These restructuring phenomena strongly depend on the components of the membranes themselves and their surroundings [19, 10, 96, 17, 73, 104, 2, 5, 98], leading to PDE systems coupling complex multiphysics to free surfaces problems. The resulting systems typically exhibit highly non-linear dynamics and a general numerical framework to simulate these rich dynamics in a unified and stable setting is still missing. The goal of this article is to present a finite element based computational framework which addresses some of the key challenges arising in continuum-mechanics based computational models of cells.
1.1 State of the art
@@
Numerous numerical methods have been proposed to deal with the challenges posed by a moving deformable membrane [61, 92, 1, 71, 93, 69, 90, 85, 7], including: immersed boundary methods, level-set method, mesh-free methods, particle methods, and the phase-field method. Despite numerous successes, even the most advanced schemes seem to lack the following features:
  1. P.1
@@
    It is not rare, in the targeted applications, to incur advection-dominant (AD) equations possibly posed on a moving surface. Characteristic examples can be the transport of surfactants in two-phase flow or the interaction between actin’s barbed ends and the cellular membrane. There exists a large bulk of literature for parabolic advection-diffusion-reaction (ADR) PDEs on moving surfaces [38, 40, 18, 80, 39, 67, 68, 42, 25]. For what concerns AD equations, the bulk setting has been extensively explored and the surface case has seen recent developments  [91, 22, 97, 23, 33, 100, 101]. Surprisingly, most aforementioned approaches only consider the case of stationary surfaces, leaving the AD case largely unaddressed. A notable exception is the work in [53], where characteristic cut finite element methods on moving surfaces is proposed.
  2. P.2
@@
    Physical structure preservation (SP) is still overlooked by the majority of the available articles. It is crucial, in order to maintain physical interpretability as well as numerical stability, to be able to guarantee constraints such as: mass and density positivity, phase bounds, energy decay, mass preservation, etc. A vast and diversified literature exist dedicated to preserve bounds, see [79, 74, 34, 75, 103, 102, 76, 28, 58, 56] and references therein. Unfortunately, this richness of options is lacking if we consider PDEs on surfaces, let alone moving surfaces.
  3. P.3
    Most articles assume idealized geometries, in which the complexity is still highly reduced with respect to realistic geometries resulting from imaging techniques [72]. This is especially crucial in this context since we expect the domain evolution to be dependent on the coupled equations and not known a priori. Steps are continuously being taken in this direction for what concerns geometric flows, which notoriously drive meshes towards bad conditioning, see [11, 81, 57, 36, 9, 51]. However, even in these cases, one is quite far from the structures considered in more applied articles such as [50].
The present work aims to propose a first step towards a cure of the missing pieces described above. It is meant to bridge the complexity gap between numerical accuracy and biophysical complexity in moving domains scenario. While there exist a plethora of numerical methods which each target a specific problem class or particular issue, we here employ and extend some of the most recent state-of-the-art methods and combine them into a holistic simulation framework which can handle coupled multiphysics and shape deformation problems arising in computational cell biology.
@@ 1.2 New contributions and outline
In Section 2, we provide the notation necessary for the formulation and discretization of PDEs posed on evolving domains. For purposes of generality, we keep the notation the same in the bulk and on the surface, only distinguishing between the two as needed. In Section 3, we adapt the structure-preserving scheme in [30] to moving bulk and surface problems as a solution to P.2. The model-agnostic, point-wise approach of [30] is fit to the first-order, piecewise-linear finite element framework here used. The result is a fast, generalizable and flexible scheme that can be readily adapted to multiple PDEs posed on moving bulk and surfaces. In this context, we use it to bound the PDE solution to physically-relevant values while at the same time preserving the total mass. In Section 4, we introduce a mesh redistribution algorithm based on the surface tangential motion of [36] and an Arbitrary Lagrangian Eulerian (ALE) approach. The two-step scheme in [36] is decoupled from the gradient flow dynamics and applied to general motions. This allows us to greatly alleviate the need of remeshing when simulating domains characterized by large deformations. At the same time, the scheme does not interfere with the shape evolution, maintaining accuracy. Both schemes in Section 3 and Section 4 take the form of a post-processing step and thus have the advantage of being a tunable feature of the code. In Section 5 we augment an ADR scheme for moving surfaces with a continuous interior penalty (CIP) stabilization to handle AD cases and solve P.1. The stabilization is proven to be convergent and so is its coupling with the structure-preserving algorithm of Section 3. The flexibility of the setting is further verified with an additional real-case example. In Section 6 we apply structure preservation to a Cahn-Hilliard type equation on moving surfaces. Numerical tests are presented that highlight how the algorithm in Section 3 is readily applicable to moving surfaces and true to the underlying physics. In Section 7, we focus on P.3, applying the mesh redistribution scheme of Section 4 to the parametric method of Barrett, Garke and Nürnberg [14] for Helfrich flow. The postprocessing step is extensively tested with state-of-the-art examples revealing good convergence properties and long-term stability. In Section 8, we couple the solvers in Section 5 through 7 together in a staggered approach. Convergence is tested with manufactured solutions [67] where the ADR algorithm and a simplified version of the Helfrich algorithm are coupled. In the same context a PDE system modeling tumor growth is also reproduced. The applicability of our approach to relevant cell biophysics is further explored simulating phase separation on a deformable membrane, where the Cahn-Hilliard and Helfrich schemes are coupled.
@@ 2 Model
This article explores numerical approaches to biophysical cell phenomena where the dynamics involve moving boundaries, such as cell membranes, that are assumed to have negligible thickness and thus are treated as moving hypersurfaces. As a result, we begin by outlining the essential notation required for both formulating and discretizing partial differential equations on evolving domains, with specific attention given to surface-bound PDEs [39, 15].
@@
(1) $\mathbb{P}_{M}({\boldsymbol{p}})=\mathbb{I}-{\boldsymbol{n}}_{M}({\boldsymbol{p}})\otimes{\boldsymbol{n}}_{M}({\boldsymbol{p}}),$
@@
(2) $\nabla_{M}f({\boldsymbol{p}})=\nabla\bar{f}({\boldsymbol{p}})-\mathbb{P}_{M}({\boldsymbol{p}})\nabla\bar{f}({\boldsymbol{p}}),$
@@
(3) $\Delta_{M}f({\boldsymbol{p}})=\nabla_{M}\cdot\nabla_{M}f({\boldsymbol{p}}).$
@@
(4) $\mathbb{H}({\boldsymbol{p}})=-\nabla_{M}{\boldsymbol{n}}_{M}({\boldsymbol{p}}).$
@@
(5) $\kappa({\boldsymbol{p}})=\mathrm{Tr}(\mathbb{H}({\boldsymbol{p}})).$
@@
(6) ${\boldsymbol{\Phi}}:J\times\widehat{M}\to\mathbb{R}^{d},\quad{\boldsymbol{\Phi}},{\boldsymbol{\Phi}}^{-1}\in C^{1}(J;C^{2}(\mathbb{R}^{d};\mathbb{R}^{d})),$
@@
The mapping ${\boldsymbol{\Phi}}_{t}$ can be used to define pull-back and push-forward maps of functions [4, 3]
(7) ${\boldsymbol{\Phi}}_{-t}f=f\circ{\boldsymbol{\Phi}}_{t}:\widehat{M}\rightarrow\mathbb{R}\text{ for }f:M(t)\rightarrow\mathbb{R},$
@@
(8) $\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}{\boldsymbol{\Phi}}_{t}({\boldsymbol{p}})={\boldsymbol{v}}(t,{\boldsymbol{\Phi}}_{t}({\boldsymbol{p}})).$
For ${\boldsymbol{x}}\in M(t)$, the velocity $\boldsymbol{v}$} can be split into a tangential component ${\boldsymbol{v}}^{\top}(t,{\boldsymbol{x}})\in T_{{\boldsymbol{x}}}M(t)$ and a normal one ${\boldsymbol{v}}^{\perp}(t,{\boldsymbol{x}})={\boldsymbol{v}}(t,{\boldsymbol{x}})-{\boldsymbol{v}}^{\top}(t,{\boldsymbol{x}})$.
This allows us to define a normal derivative of a scalar function $f$ on $M(t)$ as
(9) $\partial^{\circ}f(t,{\boldsymbol{x}})=\partial_{t}\bar{f}(t,{\boldsymbol{x}})+{\boldsymbol{v}}^{\perp}(t,{\boldsymbol{x}})\cdot\nabla\bar{f}(t,{\boldsymbol{x}}),$
@@
(10) $\partial^{\bullet}f(t,{\boldsymbol{x}})=[\partial^{\circ}f+{\boldsymbol{v}}^{\top}\cdot\nabla_{M(t)}\bar{f}](t,{\boldsymbol{x}})=[\partial_{t}\bar{f}+{\boldsymbol{v}}\cdot\nabla\bar{f}](t,{\boldsymbol{x}}).$
@@
(11) $\partial^{\mathcal{A}}f(t,{\boldsymbol{x}})=[\partial^{\bullet}f+({\boldsymbol{v}}^{\mathcal{A}}-{\boldsymbol{v}})\nabla_{M(t)}f](t,{\boldsymbol{x}}).$
@@ 2.1 Space discretization
The geometries are discretized using piecewise linear elements, where for surfaces we follow the setup of [39]. An $m$-dimensional manifold $M$ is approximated by a triangulated $m$-dimensional domain denoted by $M_{h}$. The elements of the discretization will be distinguished based on their codimension. We suppose $M_{h}$ is composed by a collection $\mathcal{T}_{h}$ of $m$-simplices which vertices ${\boldsymbol{y}}_{i},\;i=1,\ldots,N$ lie on $M$ and such that
(12) $\mathcal{T}_{h}=\{T\},\quad M_{h}=\bigcup T.$
@@
Throughout this work, we use a first-order fitted finite element method to discretize the PDEs in space [46]. In particular, for surfaces, we employ the surface finite element method (SFEM) as presented in [39]. The corresponding finite element space is defined as
(14) $V_{h}(M_{h}):=\{v_{h}\in C(M_{h})\;|\;v_{h}\rvert_{T}\text{ is linear for all }T\in\mathcal{T}_{h}\},$
@@
(15) $V_{h0}(M_{h}):=\{v_{h}\in C(M_{h})\;|\;v_{h}\rvert_{T}\text{ is linear for all }T\in\mathcal{T}_{h}\text{ and }v_{h}|_{\partial M_{h}}=0\}.$
@@
(16) $(\cdot\,,\,\cdot)_{M_{h}}=\sum_{T\in M_{h}}(\cdot\,,\,\cdot)_{T},\quad(\cdot\,,\,\cdot)_{\mathcal{F}_{h}^{i}}=\sum_{F\in\mathcal{F}_{h}^{i}}(\cdot\,,\,\cdot)_{F},$

### v2

@@ Abstract
Cellular morphodynamics requires solving systems of coupled partial differential equations on moving bulk and surface domains, where advection-dominant transport, structure preservation, and severe mesh distortions make robust simulation difficult. We present a holistic finite element framework that jointly addresses these obstacles for biophysical applications by combining model-agnostic structure-preserving postprocessing, ALE-based mesh redistribution strategies driven by surface-tangential velocities, and stabilized discretization for advection-diffusion-reaction problems tailored to evolving domains. The methodology is modular and applies to advection-diffusion-reaction systems, Cahn–Hilliard phase separation, Helfrich-type geometric flows, as well as their staggered and potentially mixed-dimensional couplings. We provide a concise notation for evolving bulk and surface geometries, extend positivity-, bound-, and mass-preserving projections to moving meshes, and develop a two-step redistribution procedure that maintains element quality without remeshing. Convergence studies, manufactured solutions, and biologically motivated test cases —including tumor-growth surrogates and phase segregation on deformable membranes— demonstrate accuracy, stability, and versatility across the problem classes considered.
@@
Biophysics , moving boundary problems , structure preservation , mesh redistribution , multiphysics coupling , bulk-surface coupling
2008 MSC:
35R37 , 74K15 , 65N30 , 76T99
journal: Computer Methods in Applied Mechanics and Engineering
[…]
{highlights}
Fast and accurate structure preserving scheme for moving bulk and surface PDEs
Flexible mesh-redistribution algorithm for biological moving boundary problems
Stabilized advection-dominant transport equations on moving surfaces
Spacetime convergence for structure preserving PDEs on moving surfaces
Application to state-of-the-art coupled problems involving deformable membranes
@@ 1.1 Background and state of the art
Many biophysical phenomena in cell biology can be characterized by the feedback between form and function. Cells routinely change their shape during cell motility, wound healing, and development. Such shape changes are a result of coordination of multiple biochemical and mechanical processes within the cell. Not surprisingly, these phenomena have been studied extensively using models based on partial differential equations (PDEs) with different levels of biophysical complexity Lomakin et al. [2015], Auddya et al. [2021]. In concert with experiments, such models have provided insight into the mechanisms underlying the change in cell shape Dayel et al. [2009], Sens [2020], Keren et al. [2008]. However, an open challenge in the field of computational cell biophysics is the development of robust computational tools for numerical solutions of the governing PDEs. For example, cellular membranes are mechanical boundaries that are characterized by lateral dimensions that are large compared to their thickness, appearing in a variety of structures in cells Steigmann [2018]. It has been proven that the properties of these surfaces are crucial for the genesis and maintenance of cellular membrane systems Baumgart et al. [2003], Mullins [2005]. These restructuring phenomena strongly depend on the components of the membranes themselves and their surroundings Brangwynne et al. [2009], Banani et al. [2017], Shin and Brangwynne [2017], Beutel et al. [2019], Lee et al. [2020b], Zhao and Zhang [2020], Alberti and Hyman [2021], Ambroggio et al. [2021], Snead et al. [2022], leading to PDE systems coupling complex multiphysics to free surfaces problems. The resulting systems typically exhibit highly non-linear dynamics and a general numerical framework to simulate these rich dynamics in a unified and stable setting is still missing. The goal of this article is to present a finite element based computational framework which addresses some of the key challenges arising in continuum-mechanics based computational models of cells.
@@
Numerous numerical methods have been proposed to deal with the challenges posed by a moving deformable membrane Iyer et al. [2023], Ong and Lai [2020], Aland et al. [2014], Laadhari et al. [2014], Rosolen et al. [2013], Krüger et al. [2011], Noguchi [2004], Mokbel et al. [2024], Bachini et al. [2023], including: parametric finite element methods, immersed boundary methods, level-set method, mesh-free methods, particle methods, and the phase-field method. We refer to Bonito and Nochetto [2020][Chapter 4-6] for a detailed review of some of these methods. However, even the most advanced schemes often do not simultaneously address the following features which are frequently crucial in the context of biophysical cell modeling.
@@
First, it is not rare, in the targeted applications, to incur advection-dominant equations possibly posed on a moving surface. Characteristic examples can be the transport of surfactants in two-phase flow Gross and Reusken [2011], Ganesan et al. [2017] or the interaction between actin’s barbed ends and the cellular membrane Bonilla-Quintana and Rangamani [2024], Doubrovinski and Kruse [2011]. There exists a large bulk of literature for parabolic advection-diffusion-reaction (ADR) PDEs on moving surfaces Dziuk and Elliott [2007], Dziuk et al. [2012], Bonito et al. [2013], Lubich et al. [2013], Dziuk and Elliott [2013], Kovács et al. [2017], Kovács and Power Guerra [2018], Elliott and Ranner [2021], Caetano et al. [2023], but primarily with a focus on the diffusion-dominant regime. For what concerns advection-dominant equations, the bulk setting has been extensively explored and the surface case has seen recent developments  Olshanskii et al. [2014], Burman et al. [2020], Simon and Tobiska [2019], Burman et al. [2019], Dedner and Madhavan [2015], Ulfsby et al. [2023], Xiao et al. [2020]. Surprisingly, most aforementioned approaches only consider the case of stationary surfaces, except for Hansbo et al. [2015], where a characteristic cut finite element method on moving surfaces is proposed.
@@
Second, the major part of the existing literature on surface PDEs pays little attention to structure preservation of physical quantities. In contrast, computational cell biophysics models often require the preservation of physical constraints —including mass conservation, density positivity, phase bounds, and energy decay— to maintain both physical interpretability and numerical stability of the computed solution. While there is a vast and diversified literature on bulk PDEs dedicated to preserve positivity, bounds and mass Lu et al. [2013], Li et al. [2020], Du et al. [2021], Li and Zhang [2020], Zhang and Shu [2011, 2010], Liu et al. [2018], Chen et al. [2019], Huang and Shen [2021], Hu and Huang [2020] this richness of options is unfortunately lacking if we consider PDEs on surfaces, let alone moving surfaces.
Finally, in the context of cell shape dynamics, the domain evolution depends on a set of coupled PDEs and is not known a priori. For classical prototypical geometric flows, such as mean curvature flow, surface diffusion flow, and Helfrich flow, a large number of numerical methods have been proposed, see, e.g. Dziuk [2008], Barrett et al. [2007], M. Elliott and Fritz [2017], Hu and Li [2022], Duan and Li [2024], Bai et al. [2024], Garcke et al. [2025], assuming mostly idealized geometries. It is well-known that for most finite-element based algorithms, such geometric flows notoriously drive surface meshes towards bad conditioning. Consequently, it is crucial to design numerical methods that can maintain good mesh quality during the shape evolution also in the presence of more realistic geometries resulting e.g. from imaging techniques Lee et al. [2020a], Yang et al. [2016], Gulati and Rudraraju [2024], or when coupled to other bulk-surface PDEs Francis et al. [2024].
@@ 1.2 New contributions and outline
The present work aims to present a unified finite-element based computational framework which focuses on the three challenges outlined above: advection-dominant transport on moving surfaces, structure-preservation, and handling of large mesh deformations in problems where multi-physics, potential multi-dimensional problems are nonlinearly coupled to the shape evolution. It is meant to bridge the complexity gap between numerical accuracy and biophysical complexity in moving domains scenario. While there exist a plethora of numerical methods which each target a specific problem class or particular issue, we here employ and extend some of the most recent state-of-the-art methods and combine them into a holistic simulation framework which can handle coupled multiphysics and shape deformation problems arising in computational cell biology.
We start in Section 2 by providing the notation necessary for the formulation and discretization of PDEs posed on evolving domains. For purposes of generality, the notation is kept general to cover both evolving surface and bulk domains, only distinguishing between the two as needed. In Section 3, we adapt the structure-preserving scheme in Cheng and Shen [2022b] to moving bulk and surface problems as a general approach to design positivity, bound and mass preserving schemes. The model-agnostic, pointwise approach of Cheng and Shen [2022b] is fit to the first-order, piecewise linear finite element framework used here. The result is a fast, generalizable and flexible approach that can be readily adapted to multiple PDEs posed on moving bulk and surfaces. In our context, it is used to bound PDE solution to physically-relevant values while at the same time preserving the total mass.
@@
Then, in Section 4, we present a mesh redistribution algorithm based on the surface tangential motion of Duan and Li [2024] and an Arbitrary Lagrangian Eulerian (ALE) approach. The two-step scheme in Duan and Li [2024] is decoupled from the gradient flow dynamics and applied to general motions. This allows us to greatly alleviate the need of remeshing when simulating domains characterized by large deformations. At the same time, the scheme does not interfere with the shape evolution, maintaining accuracy. The main advantage of the schemes presented in Section 3 and Section 4 is that they can be realized as a post-processing step and thus represent tunable features of the simulation framework. In Section 5 we augment an ADR scheme based on the evolving surface finite element method with a continuous interior penalty (CIP) stabilization to handle advection-dominant problems. We demonstrate that stabilized ADR discretization, potentially combined with a structure preservation postprocessing step, yields optimal convergent schemes for the problem class. The flexibility of the setting is further verified with an additional real-case example. Furthermore, we apply structure preservation to a Cahn-Hilliard type equation on moving surfaces in Section 6. Numerical tests are presented that highlight how the algorithm in Section 3 is readily applicable to moving surfaces and preserves the underlying physics. In Section 7, we combine the efficient mesh redistribution scheme of Section 4 with the parametric method of Barrett, Garke and Nürnberg Barrett et al. [2017a] for Helfrich flow. The postprocessing step is extensively tested with state-of-the-art examples revealing good convergence properties and long-term stability. Afterwards, in Section 8, we couple the solvers from Section 5 through 7 together in a staggered approach. Convergence is tested with manufactured solutions Kovács et al. [2017] where the ADR algorithm and a simplified version of the Helfrich algorithm are coupled. In the same context a PDE system modeling tumor growth is also reproduced. The applicability of our approach to relevant cell biophysics is further explored simulating phase separation on a deformable membrane, where the presented schemes for Cahn-Hilliard and Helfrich are coupled.
@@ 2 Model
This article explores numerical approaches to biophysical cell phenomena where the dynamics involve moving boundaries, such as cell membranes, that are assumed to have negligible thickness and thus are treated as moving hypersurfaces. As a result, we begin by outlining the essential notation required for both formulating and discretizing partial differential equations on evolving domains, with specific attention given to surface-bound PDEs Dziuk and Elliott [2013], Barrett et al. [2020].
@@
$\mathbb{P}_{M}({\boldsymbol{p}})=\mathbb{I}-{\boldsymbol{n}}_{M}({\boldsymbol{p}})\otimes{\boldsymbol{n}}_{M}({\boldsymbol{p}}),$ (1)
@@
$\nabla_{M}f({\boldsymbol{p}})=\nabla\bar{f}({\boldsymbol{p}})-\mathbb{P}_{M}({\boldsymbol{p}})\nabla\bar{f}({\boldsymbol{p}}),$ (2)
@@
$\Delta_{M}f({\boldsymbol{p}})=\nabla_{M}\cdot\nabla_{M}f({\boldsymbol{p}}).$ (3)
@@
$\mathbb{H}({\boldsymbol{p}})=-\nabla_{M}{\boldsymbol{n}}_{M}({\boldsymbol{p}}).$ (4)
@@
$\kappa({\boldsymbol{p}})=\mathrm{Tr}(\mathbb{H}({\boldsymbol{p}})).$ (5)
@@
${\boldsymbol{\Phi}}:J\times\widehat{M}\to\mathbb{R}^{d},\quad{\boldsymbol{\Phi}},{\boldsymbol{\Phi}}^{-1}\in C^{1}(J;C^{2}(\mathbb{R}^{d};\mathbb{R}^{d})),$ (6)
@@
The mapping ${\boldsymbol{\Phi}}_{t}$ can be used to define pull-back and push-forward maps of functions Alphonse et al. [2015, 2023]
${\boldsymbol{\Phi}}_{-t}f=f\circ{\boldsymbol{\Phi}}_{t}:\widehat{M}\rightarrow\mathbb{R}\text{ for }f:M(t)\rightarrow\mathbb{R},$ (7)
@@
$\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}{\boldsymbol{\Phi}}_{t}({\boldsymbol{p}})={\boldsymbol{v}}(t,{\boldsymbol{\Phi}}_{t}({\boldsymbol{p}})).$ (8)
For ${\boldsymbol{x}}\in M(t)$, the velocity $\boldsymbol{v}$} can be split into a tangential component ${\boldsymbol{v}}^{\top}(t,{\boldsymbol{x}})\in T_{{\boldsymbol{x}}}M(t)$ and a normal one ${\boldsymbol{v}}^{\perp}(t,{\boldsymbol{x}})={\boldsymbol{v}}(t,{\boldsymbol{x}})-{\boldsymbol{v}}^{\top}(t,{\boldsymbol{x}})$. This allows us to define a normal derivative of a scalar function $f$ on $M(t)$ as
$\partial^{\circ}f(t,{\boldsymbol{x}})=\partial_{t}\bar{f}(t,{\boldsymbol{x}})+{\boldsymbol{v}}^{\perp}(t,{\boldsymbol{x}})\cdot\nabla\bar{f}(t,{\boldsymbol{x}}),$ (9)
@@
$\partial^{\bullet}f(t,{\boldsymbol{x}})=[\partial^{\circ}f+{\boldsymbol{v}}^{\top}\cdot\nabla_{M(t)}\bar{f}](t,{\boldsymbol{x}})=[\partial_{t}\bar{f}+{\boldsymbol{v}}\cdot\nabla\bar{f}](t,{\boldsymbol{x}}).$ (10)
@@
$\partial^{\mathcal{A}}f(t,{\boldsymbol{x}})=[\partial^{\bullet}f+({\boldsymbol{v}}^{\mathcal{A}}-{\boldsymbol{v}})\nabla_{M(t)}f](t,{\boldsymbol{x}}).$ (11)
@@ 2.1 Space discretization
The geometries are discretized using piecewise linear elements, where for surfaces we follow the setup of Dziuk and Elliott [2013]. An $m$-dimensional manifold $M$ is approximated by a triangulated $m$-dimensional domain denoted by $M_{h}$. The elements of the discretization will be distinguished based on their codimension. We suppose $M_{h}$ is composed by a collection $\mathcal{T}_{h}$ of $m$-simplices which vertices ${\boldsymbol{y}}_{i},\;i=1,\ldots,N$ lie on $M$ and such that
$\mathcal{T}_{h}=\{T\},\quad M_{h}=\bigcup T.$ (12)
@@
Throughout this work, we use a first-order fitted finite element method to discretize the PDEs in space Ern and Guermond [2021b]. In particular, for surfaces, we employ the surface finite element method (SFEM) as presented in Dziuk and Elliott [2013]. The corresponding finite element space is defined as
$V_{h}(M_{h}):=\{v_{h}\in C(M_{h})\;|\;v_{h}\rvert_{T}\text{ is linear for all }T\in\mathcal{T}_{h}\},$ (14)
@@
$V_{h0}(M_{h}):=\{v_{h}\in C(M_{h})\;|\;v_{h}\rvert_{T}\text{ is linear for all }T\in\mathcal{T}_{h}\text{ and }v_{h}|_{\partial M_{h}}=0\}.$ (15)
@@
$(\cdot\,,\,\cdot)_{M_{h}}=\sum_{T\in M_{h}}(\cdot\,,\,\cdot)_{T},\quad(\cdot\,,\,\cdot)_{\mathcal{F}_{h}^{i}}=\sum_{F\in\mathcal{F}_{h}^{i}}(\cdot\,,\,\cdot)_{F},$ (16)
@@
Remark 1.
Together with the fitted finite element discretization explained above, an unfitted approach can also be adopted. We refer to Burman et al. [2015], Palzhanov et al. [2021], Wang et al. [2022], Olshanskii et al. [2022], Neiva and Turlier [2025], Garg and Olshanskii [2025] for advances in this direction.